### v1

@@ Abstract
Motivation: Low-complexity (LC) DNA sequences are compositionally repetitive sequences that are often associated with increased variant density and variant calling artifacts. While algorithms for identifying LC sequences exist, they either lack rigorous mathematical foundation or are inefficient with long context windows.
@@ 1 Introduction
In computer science, a string is of low complexity (LC) if it is repetitive in composition. LC strings tend to be tandemly repetitive and most of them can be identified with tandem repeat finding algorithms such as TRF (Benson, , 1999), TANTAN (Frith, , 2011), ULTRA (Olson and Wheeler, , 2024) and pytrf (Du et al., , 2025). These algorithms do not rigorously define string complexity. They rely on heuristics to search for impure tandem repeats and cannot identify LC strings without clear tandem structures. SDUST (Morgulis et al., , 2006) is the only widely used algorithm that explicitly defines string complexity and finds the exact solutions. However, with $O(w^{3}L)$ time complexity, where $w$ is the window size and $L$ is the genome length, SDUST is inefficient given a large $w$ and is thus impractical for finding satellite or tandem repeats with long motifs. Furthermore, the complexity scoring function used by SDUST is not backed by rigorous modeling. The theoretical properties of LC strings defined this way are unclear.
Inspired by SDUST, we sought an alternative way to define the $k$-mer complexity of a string and to identify LC regions in a genome. Our complexity scoring function is based on a statistical model of $k$-mer count distribution and our algorithm is practically close to $O(wL)$ in time complexity, enabling the efficient identification of LC strings in long context windows.
@@ 2 Methods
Similar to SDUST (Morgulis et al., , 2006), we define the complexity of a DNA string with a function of the $k$-mer counts of the string. In this section, we will first model the $k$-mer count distribution of random strings. We will then describe the complexity scoring function and the condition on bounding LC substrings in a long string. We will compare our method to SDUST in the end.
@@ 2.1 Notations
Let $\Sigma=\{{\tt A},{\tt C},{\tt G},{\tt T}\}$ be the DNA alphabet, $x\in\Sigma^{*}$ is a DNA string and $|x|$ is its length. $t\in\Sigma^{k}$ is a $k$-mer. For $|x|\geq k$, $c_{x}(t)$ is the occurrence of $k$-mer $t$ in $x$; $\ell(x)=\sum_{t}c_{x}(t)=|x|-k+1$ is the total number of $k$-mers in $x$. $\vec{c}_{x}$ denotes the count array over all $k$-mers.
In this article, we assume there is one long genome string of length $L$. We use closed interval $[i,j]$ to denote the substring starting at $i$ and ending at $j$, including the end points $i$ and $j$. We may use “interval” and “subsequence” interchangeably.
@@ 2.2 Modeling $k$-mer counts
Suppose symbols in $\Sigma$ all occur at equal frequency. Then for all $k$-mer $t$, $c_{x}(t)\sim{\rm Poisson}(\lambda)$ where $\lambda=\ell(x)/4^{k}$. Let
$p(n|\lambda)\triangleq\frac{\lambda^{n}}{n!}e^{-\lambda}$
be the probability mass function of Poisson distribution. Notably, although $c_{x}(t)\leq\ell(x)$, given that $\ell(x)\gg 1$ in practice,
$p(\ell|\lambda)\approx\frac{e^{-\lambda}}{\sqrt{2\pi\ell}}\cdot\left(\frac{e}{4^{k}}\right)^{\ell}\ll 1$
with the Sterling formula – $p(\ell|\lambda)$ is very close to 0. This suggests Poisson remains a good approximation.
The composite probability of string $x$ can be modeled by
$P(\vec{c}_{x})=\prod_{t\in\Sigma^{k}}p(c_{x}(t)|\lambda)$
We have
$\log P(\vec{c}_{x})=4^{k}\lambda(\log\lambda-1)-\sum_{t}\log c_{x}(t)!$ (1)
To get an intuition about $P(\vec{c}_{x})$, suppose $\ell(x)\ll 4^{k}$. In this case, $c_{x}(t)$ will be mostly 0 or 1 for a random string and the last term in Eq. (1) will be close to 0. Given an LC string of the same length, we will see more $c_{x}(t)$ of 2 or higher, which will reduce $\log P(\vec{c}_{x})$. Thus the probability of an LC string is lower under this model.
Although $\log P(\vec{c}_{x})$ can be used to compare the complexity of strings of the same length, it does not work well for strings of different lengths because $\log P(\vec{c}_{x})$ decreases with $\ell(x)$. We would like to scale it to $Q(\vec{c}_{x})$ such that $Q$ approaches 0 given a random string. We note that on the assumption of equal base frequency, the average of $\log P(\vec{c}_{x})$ can be approximated to
$H(\lambda)$ $\triangleq$ $\sum_{\vec{c}}P(\vec{c})\cdot\sum_{t}\log p(c_{t}|\lambda)$
$=$ $4^{k}\sum_{n=0}^{\infty}p(n|\lambda)\log p(n|\lambda)$
$=$ $4^{k}\lambda(\log\lambda-1)-4^{k}e^{-\lambda}\sum_{n=0}^{\infty}\log n!\cdot\frac{\lambda^{n}}{n!}$
which is the negative entropy of $P$. We can thus define
$Q(\vec{c}_{x})\triangleq H(\lambda)-\log P(\vec{c}_{x})=\sum_{t}\log c_{x}(t)!-f\left(\frac{\ell(x)}{4^{k}}\right)$
@@
$f(\lambda)\triangleq 4^{k}e^{-\lambda}\sum_{n=0}^{\infty}\log n!\cdot\frac{\lambda^{n}}{n!}$
$Q(\vec{c}_{x})$ is higher for LC string $x$.
@@ 2.3 Scoring low-complexity intervals
$S(\vec{c}_{x})\triangleq Q(\vec{c}_{x})-T\cdot\ell(x)=\sum_{t}\log c_{x}(t)!-T\cdot\ell(x)-f\left(\frac{\ell(x)}{4^{k}}\right)$ (2)
Threshold $T$ controls the level of complexity in the output. It defaults to $0.6$, less than $\log 2$. If $S(\vec{c}_{x})>0$, $x$ is considered to contain an LC substring. Note that we often do not want to classify the entire $x$ as an LC substring in this case because the concatenation of a highly repetitive sequence and a random sequence could still lead to a positive score.
Recall that we may use close intervals to represent substrings. For convenience, we write $S(\vec{c}_{[i,j]})$ as $S(i,j)$. In implementation, we precalculate $f(\ell/4^{k})$ and introduce
@@ 2.4 Finding low-complexity regions
Function FindStart() finds the longest good LC interval ending at one position. We apply the function to every position in the genome to find all good LC intervals. We can skip $j$ if [jk+1,j][j-k+1,j] is unique in [jw+1,j][j-w+1,j] because the forward pass would not reach $j$ in this case. We also introduce a heuristic to extend a good LC interval [jw,j1][j-w,j-1] to [jw+1,j][j-w+1,j] without calling FindStart(). We additionally use an X-drop heuristic (Altschul et al., , 1997) to avoid connecting two good LC intervals occasionally.
The overall longdust algorithm is inexact and may result in slightly different LC regions (21kb out of 278Mb in T2T-CHM13). We run the algorithm on both the forward and the reverse strand of the input sequences and merge the resulting intervals. The default longdust output is strand symmetric.
2.5 Comparison to SDUST
@@ 2.6 Scoring with Shannon entropy
We adapted longdust for $S_{E}$ and found using $S_{E}$ is more than twice as slow. We suspected some longdust heuristics did not work well with $S_{E}$, but we did not investigate further.
@@ 3.1 Low-complexity regions in T2T-CHM13
SDUST 4m15s 0.23 -t30
pytrf 2h39m 0.70 -M500
TRF 12h52m 7.49 2 7 7 80 10 50 500 -l12
TANTAN 32m50s 1.28 -w500 -s.85
ULTRA 146h4m 33.31 -p500 -t16
@@
Figure 1: Lengths of low-complexity regions. Low-complexity (LC) regions identified by each tool are first intersected with centromeric satellite annotation. The remainder is then intersected with longdust, TRF and SDUST in order. There are no overlaps between stacks. The total height is the length of LC regions found by each tool. Alternative settings – “TRF-c4”: requiring $\geq 4$ copies of the repeat unit; “ULTRA-s30”: requiring score $\geq 30$ in ULTRA output; “TANTAN-95”: run with “-s.95” for more stringent output.
We applied longdust, SDUST v0.1 (Morgulis et al., , 2006), pytrf v1.4.2 (Du et al., , 2025), TRF v4.10 (Benson, , 1999), TANTAN v51 (Frith, , 2011), and ULTRA v1.20 (Olson and Wheeler, , 2024) to the T2T-CHM13 human genome (Nurk et al., , 2022). Command lines for TRF, TANTAN and ULTRA were adopted from Olson and Wheeler, (2024), with the maximum period set to 500 (Table 1). Notably, TRF would not finish in days with the default option -l. An even larger -l helps performance at the cost of memory.
Longdust finds 277.1Mb of LC regions with 224.3Mb overlapping with centromeric satellite annotated by the telomere-to-telomere (T2T) consortium (Fig. 1). Of the remaining 52.7Mb, 34.1Mb overlaps with TRF; 15.4Mb of the remainder (18.6Mb) is found by SDUST. Only 3.2Mb is left, suggesting most longdust LC regions fall in centromeres or are found by TRF or SDUST.
@@
As to other tools, ULTRA outputs the largest tandem repeat regions (Fig. 1). Nevertheless, if we raise its score threshold to 30, the total region length is similar to that of TRF. TANTAN also reports many regions not found by longdust, TRF or SDUST. Increasing its probability threshold only marginally change the relative portions. Pytrf cannot effectively identify alpha satellite with $\sim$170bp repeat units, even though it was set to find tandem repeats with unit up to 500bp. SDUST is the only other method that looks for LC regions not limited to tandem repeats. With a 64bp window size, it naturally misses tandem repeats with long units, including all alpha satellite.

### v2

@@ Abstract
Motivation: Low-complexity (LC) DNA sequences are compositionally repetitive sequences that are often associated with spurious homologous matches and variant calling artifacts. While algorithms for identifying LC sequences exist, they either do not define complexity mathematically or are inefficient with long or variable context windows.
@@ 1 Introduction
A string is of low complexity (LC) if it is repetitive in composition. LC strings may lead to spurious homologous matches (Morgulis et al., , 2006) and cause variant calling artifacts (Li, , 2014; Qin and Li, , 2025). In computer science, string complexity can be measured with Kolmogorov complexity (Kolmogorov, , 1998; Silva et al., , 2022), Lempel-Ziv complexity (Lempel and Ziv, , 1976; Orlov and Potapov, , 2004) or Shannon entropy (Shannon, , 1948). In genomics, topological entropy (Crochemore and Vérin, , 1999), linguistic complexity (Troyanskaya et al., , 2002) and Markov additive process (Spouge, , 2007) have also been explored. A complication in finding LC regions is that we want to identify boundaries of local LC regions, instead of classifying an entire string as LC or not. Measuring complexity with fixed-sized sliding windows is not adequate, either, as we could not pinpoint the boundaries with such methods.
SDUST (Morgulis et al., , 2006) is a popular algorithm to identify local LC regions. It defines string complexity by running 3-mer counts and finds exact solutions under this definition. However, with the $O(w^{3}L)$ time complexity, where $w$ is the window size and $L$ is the genome length, SDUST is inefficient given a large $w$. In addition, the complexity scoring function used by SDUST grows quadratically with the string length. It tends to classify long strings as LC over short ones. A large window size $w$ would make this worse. As a result, SDUST is not suitable for finding satellite or tandem repeats with long motifs.
In genomes, LC strings are often tandemly repetitive due to DNA polymerase slippage. Such tandem repeats can be identified with tandem repeat finding algorithms such as TRF (Benson, , 1999), ULTRA (Olson and Wheeler, , 2024) and pytrf (Du et al., , 2025). They report repeat motifs and copy numbers, which are useful in downstream analyses. TANTAN (Frith, , 2011) models tandem patterns but can often identify general LC regions in practice. Unlike SDUST, these algorithms do not mathematically define string complexity.
Inspired by SDUST, we sought an alternative way to define the $k$-mer complexity of a string and to identify local LC regions in a genome. Our complexity scoring function is based on a statistical model of $k$-mer count distribution and our algorithm is practically close to $O(wL)$ in time complexity, enabling the efficient identification of LC strings in long context windows.
@@ 2 Methods
Similar to SDUST (Morgulis et al., , 2006), we define the complexity of a DNA string with a function of $k$-mer counts. In this section, we will model the $k$-mer count distribution of random strings with each $k$-mer occurring at equal frequency and describe the complexity scoring function and the condition on bounding LC substrings in a long string. We will then generalize the model to unequal $k$-mer frequency. We will also discuss alternative scoring functions such as the SDUST scoring and Shannon entropy.
@@ 2.1 Notations
Let $\Sigma=\{{\tt A},{\tt C},{\tt G},{\tt T}\}$ be the DNA alphabet, $x\in\Sigma^{*}$ is a DNA string and $|x|$ is its length. $t\in\Sigma^{k}$ is a $k$-mer. For $|x|\geq k$, $c_{x}(t)$ is the occurrence of $k$-mer $t$ in $x$; $\ell(x)=\sum_{t}c_{x}(t)=|x|-k+1$ is the total number of $k$-mers in $x$. $\vec{c}_{x}$, of size $4^{k}$, denotes the count array over all $k$-mers.
In this article, we assume there is one long genome string of length $L$. We use closed interval $[i,j]$ to denote the substring starting at $i$ and ending at $j$, including the end points $i$ and $j$. We may use “interval” and “substring” interchangeably.
@@ 2.2 Modeling $k$-mer counts
For a $k$-mer $t\in\Sigma^{k}$, $q_{t}$ is its frequency with $\sum_{t}q_{t}=1$. $\vec{c}_{x}$ follows a multinomial distribution:
$P(\vec{c}_{x};\vec{q})=\ell(x)!\prod_{t\in\Sigma^{k}}\frac{q_{t}^{c_{x}(t)}}{c_{x}(t)!}$
For convenience, introduce $r_{t}\triangleq 4^{k}q_{t}$. We have
$\log P(\vec{c}_{x};\vec{q})=g(\ell(x))-\sum_{t}\log c_{x}(t)!+\sum_{t}c_{x}(t)\log r_{t}$ (1)
@@
$g(\ell)=\log\ell!-k\ell\log 4$
is a function of string length but does not depend on $k$-mer counts.
To get an intuition about the effect of low-complexity strings, suppose $q_{t}=1/4^{k}$ and $\ell(x)\ll 4^{k}$. In this case, the last term in Eq. (1) is 0 and $c_{x}(t)$ is mostly 0 or 1 for a random string. $\log P(\vec{c}_{x})$ will be close $g(\ell(x))$. Given an LC string of the same length, we will see more $c_{x}(t)$ of 2 or higher, which will reduce $\log P(\vec{c}_{x})$. Thus the probability of an LC string is lower under this model.
Although $\log P(\vec{c}_{x})$ can be used to compare the complexity of strings of the same length, it does not work well for strings of different lengths because $\log P(\vec{c}_{x})$ decreases with $\ell(x)$. We would like to scale it to $Q(\vec{c}_{x})$ such that $Q$ approaches 0 given a random string. We note that the mean of $\log P(\vec{c}_{x})$
$H(\ell)=\sum_{\vec{c}_{x}}P(\vec{c}_{x})\log P(\vec{c}_{x})$
is the negative entropy of $P(\vec{c}_{x})$, which is
$H(\ell)=\log\ell!+\ell\sum_{t}q_{t}\log q_{t}-f(\ell)=g(\ell)+\ell\sum_{t}q_{t}\log r_{t}-f(\ell)$
where
$f(\ell;\vec{q})\triangleq\sum_{t}\sum_{n=0}^{\ell}\binom{\ell}{n}q_{t}^{n}(1-q_{t})^{\ell-n}\log n!$
Under the Poisson approximation,
$f(\ell;\vec{q})\approx\sum_{t}e^{-\ell q_{t}}\sum_{n=0}^{\infty}\log n!\cdot\frac{(\ell q_{t})^{n}}{n!}$ (2)
Now introduce
$\tilde{Q}(\vec{c}_{x})$ $\triangleq$ $H(\ell)-\log P(\vec{c}_{x})$
$=$ $\sum_{t}\log c_{x}(t)!-f(\ell(x))+\sum_{t}\big[\ell(x)q_{t}-c_{x}(t)\big]\log r_{t}$
$\tilde{Q}(\vec{c}_{x})$ approaches 0 given a random string $x$ regardless of its length. We can thus compare strings of different lengths. However, $\tilde{Q}$ measures both $k$-mer repetitiveness in the first term and $k$-mer usage in the last term. It is possible for $\tilde{Q}$ to reach a large value if string $x$ is composed of rare unique $k$-mers. To cancel the effect of rare $k$-mers, we ignore the last term and use
$Q(\vec{c}_{x};\vec{q})\triangleq\sum_{t}\log c_{x}(t)!-f(\ell(x);\vec{q})$
to measure the complexity of string $x$. $Q(\vec{c}_{x})$ also approaches 0 given a random string $x$ because $\tilde{Q}(\vec{c}_{x})$ approaches 0 and $c_{x}(t)$ approaches $\ell(x)q_{t}$. $Q(\vec{c}_{x})$ is higher for LC strings.
Figure 1: Scaling function. (a) Function value of $f(\ell)$ in Eq.(2) at $k=7$. (b) First derivative. Equivalent to per-base contribution.
In practice, $k$-mer frequency $\vec{q}$ is computed from genome-wide GC content. Low or high GC content leads to elevated $f(\ell)$ (Fig. 1), which is expected as $c_{x}(t)$ grows faster with more biased GC content.
@@ 2.3 Scoring low-complexity intervals
$S(\vec{c}_{x})\triangleq Q(\vec{c}_{x})-T\cdot\ell(x)=\sum_{t}\log c_{x}(t)!-T\cdot\ell(x)-f(\ell(x))$ (3)
Threshold $T$ controls the level of complexity in the output. It defaults to $0.6$, less than $\log 2$. At $k=7$, $f(\ell)$ is close to 0 for small $\ell$. $0.6\ell$ is the larger term within the default window size of 5kb (Fig. 1).
If $S(\vec{c}_{x})>0$, $x$ is considered to contain an LC substring. However, we often do not want to classify the entire $x$ as an LC substring in this case because the concatenation of a highly repetitive sequence and a random sequence could still lead to a positive score.
Recall that we may use close intervals to represent substrings. For convenience, we write $S(\vec{c}_{[i,j]})$ as $S(i,j)$. In implementation, we precalculate $f(\ell)$ and introduce
@@ 2.4 Finding low-complexity regions
Function FindStart() finds the longest good LC interval ending at one position. We apply the function to every position in the genome to find all good LC intervals. We can skip $j$ if [jk+1,j][j-k+1,j] is unique in [jw+1,j][j-w+1,j] because the forward pass would not reach $j$ in this case. We also introduce a heuristic to extend a good LC interval [jw,j1][j-w,j-1] to [jw+1,j][j-w+1,j] without calling FindStart().
Our definition may classify a short non-repetitive interval to LC if it is surrounded by highly repetitive LC intervals. We additionally use an X-drop heuristic (Altschul et al., , 1997) to alleviate this issue. On the T2T-CHM13 genome, the total length of LC intervals is 0.5% shorter with X-drop. The heuristic has a minor effect.
The overall longdust algorithm is inexact and may result in slightly different LC regions on opposite strands. We run the algorithm on both the forward and the reverse strand of the input sequences and merge the resulting intervals. The default longdust output is strand symmetric.
2.5 The SDUST scoring
@@ 2.6 Scoring with Shannon entropy
We adapted longdust for $S_{E}$ and found using $S_{E}$ is more than twice as slow. We suspected some longdust heuristics did not work well with $S_{E}$. Not backed up by a statistical model, it is not apparent how to integrate GC adjustment, either.
@@ 3.1 Low-complexity regions in T2T-CHM13
SDUST 0.07 0.23 -t30
pytrf 2.64 0.70 -M500
TRF 12.83 7.49 2 7 7 80 10 50 500 -l12
TANTAN 0.56 1.28 -w500
ULTRA 146.07 33.31 -p 500 -t 16
@@
We applied longdust, SDUST v0.1 (Morgulis et al., , 2006), pytrf v1.4.2 (Du et al., , 2025), TRF v4.10 (Benson, , 1999), TANTAN v51 (Frith, , 2011), and ULTRA v1.20 (Olson and Wheeler, , 2024) to the T2T-CHM13 human genome (Nurk et al., , 2022). The maximum period was set to 500 (Table 1). Notably, TRF would not finish in days with the default option -l. With -l30, TRF could run in 11 hours using 17.82GB memory at the peak.
Figure 2: Lengths of low-complexity regions in T2T-CHM13. Low-complexity (LC) regions identified by each tool are first intersected with centromeric satellite annotation. The remainder is then intersected with longdust, TRF and SDUST in order. There are no overlaps between stacks. The total height is the length of LC regions found by each tool. Alternative settings – “longdust-w8k”: window size $w=8000$; “longdust-t.5”: score threshold $T=0.5$; “longdust-k8”: using 8-mers; “longdust-g.3”: genome-wide GC set to 30%; “TRF-c4”: requiring $\geq 4$ copies of the repeat unit; “TANTAN-r1e-9”: run with “-r 1e-9”; “ULTRA-s30”: requiring score $\geq 30$ in the output.
Due to the lack of ground truth, we could not calculate the accuracy of longdust. We instead checked how often longdust reports regions unique to itself and how often it misses regions reported by other tools, especially by two or more of them.
Longdust finds 277.1Mb of LC regions with 224.3Mb overlapping with centromeric satellite annotated by the telomere-to-telomere (T2T) consortium (Fig. 2). Of the remaining 52.7Mb, 34.1Mb overlaps with TRF; 15.4Mb of the remainder (18.6Mb) is found by SDUST. Only 3.2Mb is left, suggesting most longdust LC regions fall in centromeres or are found by TRF or SDUST. Longdust uses three parameters: $k$-mer size, window size $w$ and score threshold $T$. Varying these parameters do not greatly alter the results. The GC adjustment also has minor effect even when we intentionally set the genome-wide GC below the actual value of 41%.
@@
ULTRA outputs 68.2Mb of regions not reported by longdust, TRF or SDUST. Nevertheless, if we raise its score threshold to 30, the total region length is similar to that of TRF with most of repeats of $\geq$4 copies captured by longdust. Under the default setting, TANTAN finds the largest LC regions missed by longdust, TRF or SDUST. Decreasing the repeat starting threshold from 0.005 to $10^{-9}$ greatly reduces its unique regions from 148.5Mb to 24.5Mb. We suspect the remaining regions unique to TANTAN may have low copy number but we cannot control its output. TANTAN can optionally generate repeat units but this mode is six times as slow and reports only 27% of regions in the default mode.
As to other tools, pytrf cannot effectively identify alpha satellite with $\sim$170bp repeat units, even though it was set to find tandem repeats with unit up to 500bp. With a 64bp window size, SDUST naturally misses tandem repeats with long units, including all alpha satellite.
3.2 Behaviors in long satellite repeats
Although longdust, TRF, TANTAN and ULTRA report similar amount of human satellites, the LC block sizes vary greatly. We will take the centromeric satellite of chromosome 11 for an example. Around this centromere, longdust reports a continuous block of 3.5Mb from 50,999,020 to 54,488,059 on T2T-CHM13. TRF finds a nearly identical region, with the ending position shifted by only 3bp. In contrast, although ULTRA finds 99.86% of this region, it breaks them into many smaller blocks. The longest block is 120kb in size. TANTAN under the default setting misses 2.3% of this long alpha array and its output is more fragmented. The longest block is only 1587bp. With “-r 1e-9”, TANTAN misses 7.6% of this satellite. Across all satellite regions annotated on T2T-CHM13, the N50 block size is 2.9Mb, 2.0Mb, 10.2kb and 340bp for longdust, TRF, ULTRA and TANTAN, respectively. Longdust reports the longest blocks probably because it merges satellites of different types.
One application of LC masking is to speed up cross-species whole-genome alignment. 90.8% of satellites found by longdust lie in blocks longer than 10kb and conversely, 99.2% of blocks longer than 10kb fall in annotated satellites. By setting a block length threshold, we can mask the bulk of satellite without affecting the rest of the genome. A similar strategy works for TRF that also reports long blocks. We would need to stitch shorter blocks reported ULTRA and TANTAN to use this strategy, which would add another layer of complexity.
3.3 Checking strand symmetry
Figure 3: Strand asymmetry. Each tool is run on the original T2T-CHM13 and the reverse complement of the genome. The LC intervals outputted from the two runs are compared. The Y axis shows $\rm 1-intersection/union$} for each tool. Longdust by default applies the algorithm to both strands and outputs the union; “longdust-f” refers to the mode that processes one strand only.
SDUST is strand symmetric in that running SDUST on either strand of the input sequences gives identical results. While longdust defines string complexity with strand symmetry, it uses heuristics in implementation and may report different LC intervals if it is run on the reverse complement of the input sequences. To achieve strand symmetry, longdust applies the same algorithm to both strands and output the union intervals. Without the union, 0.17% of LC regions are specific to one strand (Fig. 3).
We also checked the strand differences of other tools. For most of them, a few percent of LC regions are strand specific. However, a large fraction of LC regions found by TANTAN was only reported on one strand. The difference is driven by regions outside centromeric satellites, where the percent difference reaches 60.0%.